OpenStudy (anonymous):

A woman works at a law firm in city A which is about 65 miles from city B. She must go to the law library in city B to get a document. Find how long it takes her to drive round-trip if she averages 50 mph .

OpenStudy (triciaal):

distance d = velocity, v * time, t d = vt how would you find t? distance is round trip 65 miles times 2 = 130 miles velocity is the average speed = 50 mph find the time by putting in the known values in the equation and solving for t

OpenStudy (triciaal):

another way to look at is 50 times 1mile travelled in one hour divide by 50 means 1 mile travelled in 1/50 hour 65 miles = 65*1mile so it would take 65*1/50 hour this is one way so the round-trip would take twice the time for 1-way = 2*65*1/50 and the time is in hours

OpenStudy (triciaal):

any questions?

OpenStudy (mathstudent55):

Since the problem mentions an average speed, it's a good idea to start with an equation for the average speed. \(average~speed = \dfrac{distance}{time}\) We can use the variables s = average speed, d = distance, and t = time to make it simpler: \(s = \dfrac{d}{t}\) The problem is asking for the time, so we solve out equation for t, time: Multiply both sides by t: \(st = d\) Divide both sides by s: \(t = \dfrac{d}{s} \) Now we know that the time is the distance divided by the average speed. We need a distance and an average speed to find the time. The cities are 65 miles apart, so a round trip is \(2 \times 65 ~miles = 130 ~miles\). That's our distance. The average speed is given as 50 mph. \(t = \dfrac{d}{s} \) \(\large t = \dfrac{130~miles}{50 \frac{miles}{hour} } \) \(t = 2.6 ~miles \times \dfrac{hour}{miles} \) \( t = 2.6 ~hours\) As you can see, the units cancel out to give us hours, which is a correct unit for time.

OpenStudy (anonymous):

Thank you, to the both of you guys. I appreciate it.

OpenStudy (triciaal):

You are welcome